OpenStudy (anonymous):

Simplify: | −6 + 8i | Question 9 options: √196 10 14 + i 16 - i

OpenStudy (anonymous):

@dtan5457 i realy do not understand this

OpenStudy (dtan5457):

o i forgot this give me a second

OpenStudy (anonymous):

okay :)

OpenStudy (dtan5457):

GOT IT

OpenStudy (anonymous):

yay :)

OpenStudy (dtan5457):

what u looking at over here is the formula of \[\left| a+bi \right|=\sqrt{a^2+b^2}\]

OpenStudy (dtan5457):

so a=-6 and b=8

OpenStudy (dtan5457):

\[\sqrt{36+64}=100 \]

OpenStudy (dtan5457):

square root of 100 is...

OpenStudy (dtan5457):

10!

OpenStudy (anonymous):

10

OpenStudy (dtan5457):

exactly good job

OpenStudy (anonymous):

yes thank you but what if its a i as a b ?

OpenStudy (anonymous):

so i do i^2?

OpenStudy (anonymous):

like for example -11 + i

OpenStudy (dtan5457):

oh. when there are absolute values, you ignore the i

OpenStudy (dtan5457):

also the equation is a+bi so im pretty sure there must be a coefficient next to i.

OpenStudy (anonymous):

no there is no coefiicient

OpenStudy (dtan5457):

do you have a specific example thats from your book or something? a=always never i bi= must have coefficient or what else will b be

OpenStudy (anonymous):

Simplify: | −11 + i | Question 8 options: √122 12 11 + i 11 - i

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

oooh. simple let me show you

OpenStudy (anonymous):

okay :)

OpenStudy (dtan5457):

i still represents \[\sqrt{-1}\]

OpenStudy (dtan5457):

so therefore is you follow the formula of |a+bi|=a2+b2−−−−−−√ a which is -11 in this case will become 121 and the square root of -1 times the square root of -1 is just -1

OpenStudy (dtan5457):

so to sum it all up you get \[\sqrt{121-1}\]

OpenStudy (dtan5457):

-1 becomes i +11

OpenStudy (dtan5457):

from the 121

OpenStudy (dtan5457):

so your answer is simply 11+i

OpenStudy (dtan5457):

clear?

OpenStudy (anonymous):

yes very clear

OpenStudy (dtan5457):

nice

OpenStudy (dtan5457):

tag me if you got more questions